OpenStudy (anonymous):

Find h′(1) if h=f∘g, f(x)=−(3x^2)−7 and the equation of the tangent line of g at 1 is y=9x−8.

OpenStudy (loser66):

@e.mccormick

OpenStudy (e.mccormick):

Do you know the chain rule for composition of functions?

OpenStudy (anonymous):

mostly

OpenStudy (e.mccormick):

The \(h=f\circ g\) means you have a composition. So I am thinking you need to chain rule that, then do the evaluation at 1.

OpenStudy (anonymous):

So I did the product rule and got 6x(9x-8)+((3x^2)-7)(9) What should I do from there - because what I did after got me the wrong answer =/

OpenStudy (e.mccormick):

chain rule is this \(f(g(x))=f'(g(x))g'(x)\)

OpenStudy (e.mccormick):

hmm.... but they never actually give you g. Just the equation of the tangent line there.

OpenStudy (e.mccormick):

Hmm... have you done integrals or antiderivatives at this point?

OpenStudy (anonymous):

Nope that's later apparently for my class

OpenStudy (e.mccormick):

It is latter most of the time. I think all the g stuff will go to 1, based on what I see there.

OpenStudy (anonymous):

So what would you suggest the answer be?

OpenStudy (e.mccormick):

I think if you just work on the derivative of −(3x^2) it should get you the right answer. The -7 goes away at that point. So derive that, then evaluate it at 1.

OpenStudy (jkristia):

Since they are asking for h'(1) and you are given g(1) I think f(1) = -(3(9x-8)^2) - 7 then f'(1) = -3(2)(9x-8)(9) f'(1) = -6(1)(9) f'(1) = -54

OpenStudy (e.mccormick):

@jkristia You are not given g(1). You are given g'(1)'s equation of the tangent line.

OpenStudy (jkristia):

oh - right, my mistake

OpenStudy (e.mccormick):

Yaah, well... I was making that mistake at the start too. LOL. And since this is pre-integral work, I think they went for something simple.

OpenStudy (anonymous):

So is there any other way how to get -54 then? because that is apparently the right answer

OpenStudy (jkristia):

I don't see how to get the composite function without taking the anti-derivative of g'(1) then No actually it still works, you take derivate of f(x) - then insert g'(1), argh, no, you are missing the '9' then ... hmmm f(1) = -(3(x)^2) - 7 f'(1) = -3(2)(x) now insert g' instead of x f'(1) = -3(2)(9x-8)(9) <-- missing '9' f'(1) = -6(1)(9) f'(1) = -54

OpenStudy (e.mccormick):

Yah... with it pre-integration I was a tad stumped on that. LOL.

OpenStudy (e.mccormick):

Wait... what if the fact that g is a tangent line is incidental!

OpenStudy (e.mccormick):

.... no.... "tangent line OF g" so not incidental...

OpenStudy (jkristia):

Hmm, I thought I had a solution, but didn't work.. I would be interested in knowing the correct solution (not just result) for this

OpenStudy (e.mccormick):

Yah... same boat as me. It seems... odd... I am sure it is either something simple we missed, or a poorly written question, with no middle ground!

OpenStudy (jkristia):

most likely something we (I) missed - it happens a lot to me haha

OpenStudy (e.mccormick):

The g'(1)=9 doh! And I even thought about that before I got lost trying to find g()

OpenStudy (dan815):

are u done this question or u want solution?

OpenStudy (e.mccormick):

We found the solution and our mistake in finding it. LOL @Invizen Did you get the point from that link to a related problem, or still need more?

OpenStudy (dan815):

ok :)

OpenStudy (anonymous):

I'm all good now thanks so much!

OpenStudy (jkristia):

>> round about way of telling you that g ′ (3)=−3 ahhh, that Is what I was missing, so in this case So g(1) = 9x-8 = 1 g'(1) = 9 f'(x) = -6x h'(1) = f'(g(1))g'(1) h'(1) = -6(1)(9) h'(1) = -54

OpenStudy (e.mccormick):

Yah, "equation of"... slope... LOL. y=mx+b came and bit us on the anatomy!